# The Broader Landscape of Robustness in Algorithmic Statistics

### v1

@@ 4.3 More Connections with Contamination Robustness
The above discussion shows that contamination-robust estimators imply private ones. Does the reverse hold? It is folklore that private algorithms are automatically contamination robust, with (small) contamination parameter $\eta\approx 1/\varepsilon$. This is not hard to show by a property of differential privacy known as group privacy. While differential privacy guarantees the similarity of an algorithm’s output distributions on inputs at distance $1$, the group privacy property guarantees similarity on inputs at distance $k$, albeit with a quantitative degradation in similarity by a factor of $k$. Setting $k\approx 1/\varepsilon n$, an easy calculation gives the desired result. Asi, Ullman, and Zakynthinou [AUZ23] build upon this fact to show that contamination robustness and privacy are equivalent for low-dimensional problems. Georgiev and Hopkins [GH22] further show that any private algorithm with very high success probability can enjoy even stronger contamination robustness, up to the parameter regime $\eta=\Omega(1)$. Note that this type of high probability guarantee is reminiscent of the sub-Gaussian rate we were aiming for in Section 3, hinting at even more connections.
@@ Acknowledgments
GK would like to thank Gavin Brown, Yeshwanth Cherapanamjeri, Samuel B. Hopkins, Mahbod Majid, Argyris Mouzakis, and Lydia Zakynthinou for their helpful comments while writing this article. GK is supported by a Canada CIFAR AI Chair, an NSERC Discovery Grant, and an unrestricted gift from Google.

### v2

@@ 4.3 More Connections with Contamination Robustness
The above discussion shows that contamination-robust estimators imply private ones. Does the reverse hold? It is known that private algorithms are automatically contamination robust, with (small) contamination parameter $\eta\approx 1/\varepsilon{{n}}$. This is straightforward as a consequence of an elementary property of differential privacy known as group privacy. While differential privacy guarantees the similarity of an algorithm’s output distributions on inputs at distance $1$, the group privacy property guarantees similarity on inputs at distance $k$, albeit with a quantitative degradation in similarity by a factor of $k$. Setting $k\approx 1/\varepsilon n$, an easy calculation gives the desired result. Asi, Ullman, and Zakynthinou [AUZ23] build upon this fact to show that contamination robustness and privacy are equivalent for low-dimensional problems. Georgiev and Hopkins [GH22] further show that any private algorithm with very high success probability can enjoy even stronger contamination robustness, up to the parameter regime $\eta=\Omega(1)$. Note that this type of high probability guarantee is reminiscent of the sub-Gaussian rate we were aiming for in Section 3, hinting at even more connections.
@@ Acknowledgments
GK would like to thank Gavin Brown, Yeshwanth Cherapanamjeri, Samuel B. Hopkins, Mahbod Majid, Argyris Mouzakis, Lydia Zakynthinou, and the anonymous reviewers for their helpful comments while writing this article. GK is supported by a Canada CIFAR AI Chair, an NSERC Discovery Grant, and an unrestricted gift from Google.